The following reaction was allowed to go to completion and produced a total of 30.4 mg NH3. If the energy transfers into 350 grams of air and increases the temperature only 4⁰C, what is the specific heat of air?
N2 + 3H2 --> 2NH3 ∆H = - 92.6 kJ /mol
1
Expert's answer
2020-12-10T07:47:18-0500

"N2 + 3H2 \to 2NH3\\
 ∆H = - 92.6 kJ /mol"


From the reaction above, when 2 moles of NH3 is produced, 92.6kJ is evolved


This means that 34g of NH3 = 92.6kJ

30.4mg = xkJ


x = (30.4/1000g × 92.6)/34g = 0.0828kJ = 82.8J.

This means 82.8J is evolved in the reaction.


Using, H = mc∆T

H = 82.8J

m = 350g

∆T = 4


c = H/m∆T = 82.8/350×4 = 0.0591 J/g.°C